# Design of high-strength, radiopure copper-chromium alloys for rare-event searches assisted by computational thermodynamics

### v1

@@ 4.1 Solution heat treatment
Figure 2 shows the Cr concentration profile obtained after 1D simulations using DICTRA for the two cases. Experimental measurements obtained from heat treated and aged samples, for the 3 out of 5 temperatures are included from Vitale et al. [17].
According to simulations, for the 1 $\mu m$ Cr thickness (Fig. 2(a)), a single phase can be achieved after 24 hours at 1025C, 1050C and 1060C. A homogeneous alloy composition of about 0.2 wtCr is obtained from simulations at 1050C and 1060C. In the case of 1025C, a longer duration is required to achieve homogenization, namely 72 h, as shown in Fig. 3. Thermodynamic stability is achieved faster at higher temperatures, as shown from the simulation results in Fig. 2(a) and Fig. 3.
@@
In the case of 1 $\mu m$ Cr in contact with 400 $\mu m$ of Cu, simulations of the solution heat treatment stage are found in agreement with the experimental data of the heat treated and aged samples of Ref. [17], in particular for the 950C and 1000C 24-hour isothermals. It is noted that in Ref. [17] Cr concentration profiles in the Cu region are reported at a distance larger than 25 $\mu m$ from the interface. Moreover, although measurements in Ref. [17] correspond to samples which are also aged, the observed trend is reproduced by the simulations, in particular for the lowest temperatures.
@@
From the presented comparisons, we deduce that the optimal solution heat treatment is the 24 h at 1050C, rather than the 1025C employed in Ref. [17]. In addition, the 1 $\mu m$ Cr layer in contact with 400 $\mu m$ Cu allows homogenization of the Cu-0.2Cr (in wt%) alloy. To achieve the maximum amount of Cr in Cu, in order to subsequently maximize the volume fraction of Cr precipitates during aging, the Cr layer thickness needs to be increased. Additionally, simulations show that a 10 $\mu m$-thick Cr layer in contact with a 400 $\mu m$-thick Cu layer is too thick for the Cr to be fully incorporated in the Cu within 24 h and give a homogenized alloy. The optimum layer configuration is the subject of further investigation [43].
@@ 4.2 Precipitation strengthening
Figure 4: Results of simulations of aging Cu-0.5Cr at 400C, 450C, and 500C, for 4(a) the evolution of volume percentage of precipitate, 4(b) Cr concentration in matrix, and 4(c) mean radius. Experimental data corresponding to aging at 450C are included [9].
@@
Aging proceeds slower at lower temperatures. The percentage of volume fraction of the precipitating phase is limited to 0.61% for all temperatures, reaching a plateau. The Cr concentration in the matrix decreases with time. It is 0.5 wtin the homogenized initial composition and as the precipitating phase bcc forms, Cr diffuses from the fcc matrix at almost 100% of its initial concentration. The results of this work using TCHEA6 and MOBHEA3 are very similar to simulations from literature for the 450[9] using the Thermo-calc Software Cu-based alloys databases TCCU3 and MOBCU3 [44, 45], especially in the % of volume fraction of precipitates and the Cr concentration in the matrix.
@@
In terms of mean radius, the simulations of this work at 450C predict a smaller radius of precipitates compared to the simulation results reported in Ref. [9], which explains the higher yield strength due to precipitation strengthening predicted in this work, as shown in Fig. 5(b). The final mean radius at 400C is predicted to be the same as that at 450C, although typically the maximum radius of the precipitates decreases with temperature. With smaller precipitates and the same volume fraction, even more enhanced mechanical properties are expected. This is shown in the graph with the prediction of the increase in yield strength due to precipitation strengthening (Fig. 5(b)). Interestingly, the same maximum radius at the plateu found for aging at 450C and 400C is not reflected to the yield strength due to precipitation strengthening. This can be explained by the fact that at lower temperatures precipitation growth proceeds slower and the volume fraction of the precipitating phase is still growing with time and has not yet reached its plateau, as shown in Fig. 4(a). The comparison of temporal evolution of mean radius and precipitate volume fraction becomes apparent in Eq. 4.
@@
Predictions for the evolution of yield strength due to solid solution strengthening, precipitation strengthening, and the total yield strength can be found in Figure 5 during aging at 400C, 450C, 500C. The yield strength due to solid solution strengthening data (Fig. 5(a)) is obtained by using Eq. 2 and due to the precipitation strengthening data (Fig. 5(b)) by using Eq. 3 and Eq.4. Simulations have also revealed that the maximum total yield strength, that is, the sum of solid solution and precipitation strengthening, can be achieved by aging at a lower temperature for longer periods of time, as shown in Fig. 5(c).
@@
At lower temperatures, Cr diffuses slower from the matrix toward the precipitates compared with aging at higher temperatures, and the volume fraction of Cr in the matrix is higher. This leads to an increase in the yield strength during both solid solution and precipitation strengthening. The relation is illustrated in the equations for the yield strength curves, Eq.2 and Eq.3-4, respectively. The yield strength evolution due to solid solution strengthening depends on the temporal Cr concentration in the matrix, so the trend of both is almost identical. The mole fraction of the initial concentration of Cr in the matrix for this work and Ref. [9] is identical to expected, corresponding to 0.5 wtCr. However, much lower initial yield strength has been reported in Ref. [9], with the source of this difference remains unclear. A small deviation of the results of this work compared to Ref. [9] is observed for the temporal evolution of the volume fraction of the precipitating phase, the Cr concentration in the matrix, and the size of mean radius for aging at 450C. This could be attributed to the different sets of CALPHAD-type thermodynamic and kinetic databases used in Ref. [9].
@@
From this work, we deduce that when aging at 500C is performed, much shorter durations with respect to those reported in Ref. [17] are required to obtain optimal results for precipitation strengthening. Prolonged aging at 500C leads to an increase in the mean radius of the precipitates, as some precipitates will grow larger and some others will shrink, indicating that coarsening is already taking place. Consequently, aging at lower temperatures should be further explored.
@@
Figure 6: Calculation of TTT diagram for Cu-0.5Cr for 0.01 and 0.6 volume fraction % using TCHEA6 and MOBHEA3. Experimental data is included [9].
@@
Figure 6 shows the Time-Temperature-Transformation (TTT) diagram calculated using the TCHEA6 and MOBHEA3 databases. This diagram is useful to understand the time-temperature dependence of growing a new phase, starting from a specific microstructure. According to this TTT diagram the maximum percentage of volume fraction of 0.61% can be reached after 10-12 minutes, 60 minutes and 300 minutes for aging at 500C, 450C and 400C, respectively.
@@
Moreover, the diagram is compared to the corresponding diagram in Ref. [9]. The set of High Entropy alloys databases (TCHEA6/ MOBHEA3) used in this work predict a narrower region at which bcc precipitation can occur in a wide range of aging temperatures. The observed differences are attributed to the different databased used. The calculations in Ref. [9] were performed using the Cu-based alloys databases TCCU3 and MOBCU3. These latter calculations are validated with measurements, but not yet for radiopure alloys.
@@
From the above comparisons, it is seen that any differences in the simulation results when different sets of databases are used are typically small. Further validation of databases will determine their suitability or may lead to further work developing more accurate descriptions through dedicated databases for radiopure Cu-based alloys. Already, this work validates the High Entropy alloys database set for the stage of solution heat treatment, through the presented comparisons with measurements reported in Ref. [17]. In the prediction of the temporal evolution of yield strength due to solid solution and precipitation strengthening, further comparisons with the experimental data would be beneficial. For example, the value of interfacial energy between the bcc and fcc phases is crucial, especially when the nucleation barrier is determined as shown in Eg.1. Due to the lack of an experimental value for the interfacial energy, this was taken from calculations reported in Ref. [9] using the TCCU3 and MOBCU3 databases and may affect predictions of the evolution of yield strength. This is not very easy to obtain experimentally. Further experimental measurements in yield strength will be beneficial for more accurate predictions.
@@ 5 Conclusion
Thermally processed, electroformed copper-chromium (CuCr) alloys are stronger than EFCu and have potential to maintain high radiopurity. As a result, they are crucial for the development of next-generation rare-event search experiments.
@@
Furthermore, a methodology is proposed to predict materials properties and optimize manufacturing processing, using DICTRA and TC-PRISMA simulations. This can lead to the acceleration of the design of the Cr and Cu layers configuration in the fabrication stage and the optimization of the parameters for thermal processing to achieve the maximum enhancement of mechanical properties.
Future work includes optimizing the Cu and Cr layer configuration to achieve homogeneous alloy compositions with up to 0.5 wtCr during solution heat treatment, and of the aging process aiming at lower temperatures and durations while further enhancing the alloy strength. These need to be achieved without compromising the radiopurity and electrical conductivity requirements.

### v2

@@ 4.1 Solution heat treatment
Figure 2 shows the Cr concentration profile obtained after 1D simulations using DICTRA for the two cases. Experimental measurements obtained from heat-treated and aged samples for three out of five temperatures are included from Vitale et al. [19].
According to simulations, for the 1 $\mu m$ Cr thickness (Fig. 2(a)), a single phase can be achieved after 24 hours at 1025C, 1050C and 1060C. A homogeneous alloy composition of about 0.2 wt% Cr is obtained from simulations at 1050C and 1060C. In the case of 1025C, a longer duration is required to achieve homogenization, namely 72 h, as shown in Fig. 3. Thermodynamic stability is achieved faster at higher temperatures, as shown from the simulation results in Fig. 2(a) and Fig. 3.
@@
In the case of 1 $\mu m$ Cr in contact with 400 $\mu m$ Cu, simulations of the solution heat treatment stage are found to be in agreement with the experimental data of the heat-treated and aged samples of Ref. [19], in particular for the 950C and 1000C 24-hour isothermals. It is noted that in Ref. [19] Cr concentration profiles in the Cu region are reported at a distance larger than 25 $\mu m$ from the interface. Moreover, although the measurements in Ref. [19] correspond to samples that are also aged, the observed trend is reproduced by the simulations, in particular for the lowest temperatures.
@@
From the presented comparisons, we deduce that the optimal solution heat treatment is the 24 h at 1050C, rather than the 1025C employed in Ref. [19]. In addition, the 1 $\mu m$ Cr layer in contact with 400 $\mu m$ Cu allows homogenization of the Cu-0.2Cr (in wt%) alloy. To achieve the maximum amount of Cr in Cu, in order to subsequently maximize the volume fraction of Cr precipitates during aging, the Cr layer thickness needs to be increased. Additionally, simulations show that a 10 $\mu m$-thick Cr layer in contact with a 400 $\mu m$-thick Cu layer is too thick for the Cr to be fully incorporated in the Cu within 24 h and give a homogenized alloy. Improved Cu and Cr layer configuration is the subject of further investigation [17].
@@ 4.2 Precipitation strengthening
Figure 4: Results of simulations of aging Cu-0.5Cr at 400C, 450C, and 500C, for 4(a) the evolution of volume percentage of precipitate, 4(b) Cr concentration in matrix, and 4(c) mean radius. Calculations (using Thermo-Calc Cu-based databases) corresponding to aging at 450C are included [10].
@@
Aging proceeds slower at lower temperatures. The percentage of volume fraction of the precipitating phase is limited to 0.61% for all temperatures, corresponding to the maximum allowed for the specific alloy composition, reaching a plateau. Given that in this work, the precipitation strengthening of a CuCr alloy of 0.5 wt% Cr content is explored, the volume fraction of the precipitating phase cannot exceed the particular value. The Cr concentration in the matrix decreases with time. It is 0.5 wt% in the homogenized initial composition and as the precipitating phase bcc forms, Cr diffuses from the fcc matrix at almost 100% of its initial concentration. The results of this work using TCHEA6 and MOBHEA3 are very similar to simulations from literature for the 450[10] using the Thermo-calc Software Cu-based alloys databases TCCU3 and MOBCU3 [45, 46], especially in the % of volume fraction of precipitates and the Cr concentration in the matrix.
@@
In terms of mean radius, the simulations of this work at 450C predict a smaller radius of precipitates compared to the simulation results reported in Ref. [10], which explains the higher yield strength due to precipitation strengthening predicted in this work, as shown in Fig. 5(b). The final mean radius at 400C is predicted to be the same as that at 450C, although generally the maximum radius of the precipitates decreases with temperature. With smaller precipitates and the same volume fraction, even more enhanced mechanical properties are expected. This is shown in the graph with the prediction of the increase in yield strength due to precipitation strengthening (Fig. 5(b)). Interestingly, the same maximum radius on the plateau found for aging at 450C and 400C is not reflected in the yield strength due to precipitation strengthening. This can be explained by the fact that at lower temperatures precipitation growth proceeds slower and the volume fraction of the precipitating phase is still growing with time and has not yet reached its plateau, as shown in Fig. 4(a). The comparison of the temporal evolution of the mean radius and the precipitate volume fraction becomes apparent in Eq. 4.
@@
Predictions for the evolution of the yield strength due to solid solution strengthening, precipitation strengthening, and the total yield strength can be found in Fig. 5 during aging at 400C, 450C, 500C. The yield strength (or stress) due to solid solution strengthening data (Fig. 5(a)) is obtained by using Eq. 2 and due to the precipitation strengthening data (Fig. 5(b)) by using Eq. 3 and Eq.4. Simulations have also revealed that the maximum total yield strength (or stress), that is, the sum of the corresponding data due to solid solution and precipitation strengthening, can be achieved by aging at a lower temperature for longer periods of time, as shown in Fig. 5(c).
@@
At lower temperatures, Cr diffuses slower from the matrix toward the precipitates compared with aging at higher temperatures, and the volume fraction of Cr in the matrix is higher. This leads to an increase in the yield strength during both solid solution and precipitation strengthening. The relation is illustrated in the equations for the yield strength curves, Eq.2 and Eq.3-4, respectively. The yield strength evolution due to solid solution strengthening depends on the temporal Cr concentration in the matrix, so the trend of both is almost identical. The mole fraction of the initial concentration of Cr in the matrix for this work and in Ref. [10] is identical to expected, corresponding to 0.5 wt% Cr. However, a much lower initial yield strength has been reported in Ref. [10], with the source of this difference remains unclear. A small deviation of the results of this work compared to Ref. [10] is observed for the temporal evolution of the volume fraction of the precipitating phase, the Cr concentration in the matrix, and the size of mean radius for aging at 450C. This could be attributed to the different sets of CALPHAD-type thermodynamic and kinetic databases used in Ref. [10].
@@
From this work, we deduce that when aging at 500C is performed, much shorter durations than those reported in Ref. [19] are required to obtain optimal results for precipitation strengthening. Prolonged aging at 500C leads to an increase in the mean radius of the precipitates, as some precipitates grow larger and some others will shrink, indicating that coarsening is already taking place. Consequently, aging at lower temperatures should be further explored.
@@
Figure 6: Calculation of TTT diagram for Cu-0.5Cr for 0.01 and 0.6 volume fraction % using TCHEA6 and MOBHEA3. Calculations (using Thermo-Calc Cu-based databases) are included [10].
@@
Figure 6 shows the Time-Temperature-Transformation (TTT) diagram calculated using the TCHEA6 and MOBHEA3 databases. This diagram is useful to understand the time-temperature dependence of growing a new phase, starting from a specific microstructure. According to this TTT diagram the maximum percentage of the volume fraction of 0.61% can be reached after 10-12 minutes, 60 minutes and 300 minutes for aging at 500C, 450C and 400C, respectively.
@@
Moreover, the diagram is compared to the corresponding diagram in Ref. [10]. The set of High Entropy alloys databases (TCHEA6/ MOBHEA3) used in this work predict a narrower region at which bcc precipitation can occur in a wide range of aging temperatures. The observed differences are attributed to the different databases used. The calculations in Ref. [10] were performed using the Cu-based alloys databases TCCU3 and MOBCU3. These latter calculations are validated with measurements, but not yet for radiopure alloys.
@@
From the above comparisons, it is seen that any differences in the simulation results when different sets of databases are used are typically small. Further validation of databases will determine their suitability or may lead to further work developing more accurate descriptions through dedicated databases for radiopure Cu-based alloys. Already, this work validates the High Entropy alloys database set for the stage of solution heat treatment, through the presented comparisons with measurements reported in Ref. [19]. In the prediction of the temporal evolution of yield strength due to solid solution and precipitation strengthening, further comparisons with the experimental data would be beneficial. For example, the value of interfacial energy between the bcc and fcc phases is crucial, especially when the nucleation barrier is determined as shown in Eq.1. Due to the lack of an experimental value for the interfacial energy, this was taken from calculations reported in Ref. [10] using the TCCU3 and MOBCU3 databases and may affect predictions of the evolution of yield strength. This is not very easy to obtain experimentally. Further experimental measurements in yield strength will be beneficial for more accurate predictions.
@@ 5 Conclusion
Thermally processed, electroformed copper-chromium (CuCr) alloys are stronger than EFCu and have the potential to maintain high radiopurity. As a result, they are crucial for the development of next-generation rare-event search experiments.
@@
Furthermore, a methodology is proposed to predict materials properties and optimize manufacturing processing, using DICTRA and TC-PRISMA simulations. This can lead to the acceleration of the design of the Cu and Cr layer configuration in the fabrication stage and the optimization of the parameters for thermal processing to achieve the maximum enhancement of mechanical properties.
Future work includes optimizing the Cu and Cr layer configuration to achieve homogeneous alloy compositions with up to 0.5 wt% Cr during solution heat treatment, and of the aging process aiming at lower temperatures and durations while further enhancing the alloy strength. These need to be achieved without compromising the radiopurity and electrical conductivity requirements.